How much is to be invested now at 10% compound interest p.a. if after 5 years the investment is to reach a target sum of $30000.


1
Expert's answer
2021-11-15T18:03:06-0500

"A=P(1+\frac {r} {100}) ^n"


"30,000=P(1 +0.1)^5"


"30000=(1.1)^5 P"


"1.6105P=30,000"


"P=\$18,627.76"